OpenStudy (anonymous):

area of a 4cm square

OpenStudy (anonymous):

4^2

OpenStudy (anonymous):

A=a^2 for a square.

OpenStudy (sidsiddhartha):

area of a square is=side*side u side is given 4cm so are=4*4=16

OpenStudy (anonymous):

So,\[4\times4=16 cm.^{2}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

make sure you know the fundamentals